OpenStudy (anonymous):

Solve for x. –3(5 – x) = 4x A. –2 B. 15 C. 2 D. –15

OpenStudy (anonymous):

First distribute -3. -15+3x=4x. Combine like terms: Subtract 3x -15=x x=-15 Now to check: -3(5-(-15))=4(-15) -3(20)=4(-15) -60=-60 It checks out.

OpenStudy (anonymous):

D = -15